OpenStudy (anonymous):

You save dimes and quarters in a cup in your locker. When you count them, you find that you have 20 coins with a value of $3.20. How many of each type of coin do you have? A. 10 dimes and 10 quarters B. 14 dimes and 6 quarters C. 12 dimes and 8 quarters D. 8 dimes and 12 quarters

OpenStudy (anonymous):

how much is a dime worth?

OpenStudy (anonymous):

how much is a quarter worth?

OpenStudy (anonymous):

say that we call x the number of dimes and y the number of quarters.

OpenStudy (anonymous):

is it b? @tomo

OpenStudy (anonymous):

.1x + .25y = $3.20

OpenStudy (anonymous):

x+y = 20

OpenStudy (anonymous):

solve the system of equations.

OpenStudy (anonymous):

D. thank you @tomo

jimthompson5910 (jim_thompson5910):

let d = # of dimes q = # of quarters The first equation is d+q = 20 since you have 20 coins total Solve for q to get q = 20 - d ----------------------------------------------- Total value of all coins is $3.20, which means 0.10d + 0.25q = 3.20 10d + 25q = 320 ... Multiply both sides of the last equation above by 100 to clear out the decimals. Now plug in q = 20 - d to get 10d + 25q = 320 10d + 25(20 - d) = 320 10d + 500 - 25d = 320 Solve that for d Once you have the value of d, plug it into q = 20 - d to find the value of q. I'll let you do this.